OpenStudy (anonymous):

PLEASE HELP MEDAL AND FAN I NEED HELP FAST!!! An adventure game has a number cube with 12 equal-size faces, each labeled with a number from 1 to 12. What is the probability of rolling a 4 or a multiple of 3? Express your answer in simplest form.

OpenStudy (anonymous):

please help me

OpenStudy (skullpatrol):

How many multiples of 4 are there from 1 to 12?

OpenStudy (anonymous):

3?

OpenStudy (skullpatrol):

correct :-) how many 4's are there?

OpenStudy (anonymous):

in what?

OpenStudy (skullpatrol):

A number cube with 12 equal-size faces, each labeled with a number from 1 to 12.

OpenStudy (anonymous):

3?

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

How many 4's on the wheel?

OpenStudy (anonymous):

1

OpenStudy (skullpatrol):

Correct, what is the probability a pointer will land on it?

OpenStudy (anonymous):

1/12

OpenStudy (skullpatrol):

correct

OpenStudy (anonymous):

so the answer is 1/12

OpenStudy (skullpatrol):

not yet you have to add in the multiples of 3

OpenStudy (anonymous):

ok how do we do that?

OpenStudy (skullpatrol):

what are the multiples of 3 on the wheel?

OpenStudy (anonymous):

4

OpenStudy (skullpatrol):

what are their values?

OpenStudy (anonymous):

3

OpenStudy (skullpatrol):

and?

OpenStudy (anonymous):

0.0

OpenStudy (skullpatrol):

do you know what a multiple is?

OpenStudy (anonymous):

no *sigh*

OpenStudy (skullpatrol):

The product of any real number and an integer is a multiple of the real number.

OpenStudy (skullpatrol):

$$\Huge a*n=m$$ m is the multiple n is the integer a is any real number

OpenStudy (skullpatrol):

3, 6, 9, 12 are all multiples of 3 on the wheel

OpenStudy (skullpatrol):

3*1= 3 3*2= 6 3*3= 9 3*4= 12

OpenStudy (skullpatrol):